Extending the working lifetime of the Metop-A weather satellite

Metop-A spacecraft status

 

Extending the working lifetime of the Metop-A weather satellite
Extending the working lifetime of the Metop-A weather satellite

Metop-A completed its operational mission on 15 November 2021, after achieving a lifetime in orbit of over 15 years. This was followed by the de-orbiting operations in line with debris mitigation standards. The traffic light chart and associated footnotes below represent the last status of each subsystem and instrument prior to Metop-A's removal from operational service.

Last Updated

18 November 2021

Published on

18 November 2021

Last update: 15/11/2021

Subsystem Status/Trend
Service Module AOCS   Power   DHSA  
COMMS   Housekeeping      
Thermal   PMCIF      
Payload Module PMC   TCU   PCU  
PDU   RTU   FMU  
SSR   XBS   NIU  
HRPT Off LRPT Off    
Instrument ASCAT Off MHS   ADCS Off
AMSU-A1   GRAS   SARR Off
AMSU-A2   GOME   SARP Off
HIRS   IASI      
AVHRR   SEM      

Footnotes

ADCS 2021-11-15: A-DCS off
AMSU-A1 2009-12-01: H7 declared failed.
2013-07-15: H3 NEDT out of spec (0.4 K)
2016-04-18: H8 declared failed.
ASCAT-ICU 2021-11-15: ASCAT-ICU off
ASCAT 2021-11-15: ASCAT off
COMMS 2011-06-23: Transponder A output power TM decreased
GOME 2012-09-22: throughput loss in UV stabilised with limited performance degradations
FMU 2019-08-03: FMU on B-side after PL_SOL_7
HIRS 2015-11-20: H10 out of specification limit
2016-02-23: H10 back in specification
2016-07-14: H5 and H10 out of specification limit
2016-11-23: H1, H2, H4, H5, H6, H9, and H10 out of specification limit
2017-01-30: H1 - H12 out of specification limit
2017-04-29: All channels out of specification limit apart from H19
2018-11-29: All channels out of specification limit
SEM 2019-05-17: TED Electron CDEM bias voltage at level 7, which is the maximum.
2021-08-02 Proton CDEM bias voltage at level 7, which is the maximum.
PMC 2011-10-22: PMC on B-side after PL-SOL_06
IASI 2015-04-09: IMS swap to b-side
2015-10-07: CD permanently stopped
SARR 2021-11-11: SARR Receiver and Transmitter on B-Side
2021-11-15: SARR off
SARP 2021-11-15: SARP off
MHS 2011-12-06: Local oscillator swap improved H3/4 NEDT, but trend continues
2014-03-01: H3/H4 further degrading but NEDT still within spec
2014-10-23: H3/H4 stable and improving
2016-12-06: Local oscillator swap improved H2 NEDT
2020-02-26: H2 declared failed
2021-02-22: swap to PIE-B after LO5 thermistor failure
HRPT 2011-01-18: HRPT-B unit in restricted operation. Complete longitudinal coverage zone active (No transmission at higher latitudes in both hemispheres)
2013-05-16: Switching changed from modulation to SSPA
2020-03-02: Switching changed from SSPA to modulation
2021-11-15: HRPT-B off
LRPT 2007-01-26: switched-off because of HIRS interferences